OpenStudy (anonymous):

Determine g(x) from f(x)=x^2 +5 upward 11 units?

OpenStudy (anonymous):

all you need to do is add 11 to the equation of f(x) so g(x) = x^2 + 5 + 11 g(x) = ?

OpenStudy (anonymous):

add 11 to the 5?

OpenStudy (anonymous):

Yes :) Do you know how to figure out these equations with translations?

OpenStudy (anonymous):

not a clue me and math don't get along

OpenStudy (anonymous):

How does this one look?

OpenStudy (anonymous):

do you know how to graph a parabola given the equation?

OpenStudy (anonymous):

They're asking if you can add f(x) and g(x) f(x) + g(x) can you plug f(x) = -6x -1 into the upper equation? What do you get?

OpenStudy (anonymous):

what happened to the -1?

OpenStudy (anonymous):

nvm I just looked at it and answered my own question

OpenStudy (anonymous):

What was your answer?

OpenStudy (anonymous):

ok now I'm lost

OpenStudy (anonymous):

Please stick with me! Back to where I was: f(x) + g(x) This is what they want. Do you get this?

OpenStudy (anonymous):

(f+g)(x) = f(x) + g(x) Do you understand why?

OpenStudy (anonymous):

so its b vise d?

OpenStudy (anonymous):

No, you need to work this problem out to understand . Not guessing the answer

OpenStudy (anonymous):

Trust me its better off if I guess I have a TBI you may show me then I will forget how it was done

OpenStudy (anonymous):

but if you get this once and repeatedly do it again, maybe 1/2 or even 1/8 of it sticks :o practice maynot make perfect, but better anyways, plug -6x-1 into f(x) of this equation: f(x) + g(x) you'd get -6x-1 + g(x) g(x) = 4x^2 can you plug that into g(x)? What would you get?

OpenStudy (anonymous):

plug 4x^2 into g(x) of this equation -6x-1 + g(x) I worked out more than half of this for you already.

OpenStudy (anonymous):

g(4x^2) like that?

OpenStudy (anonymous):

I mean g(x) = 4x^2 so replace g(x) with 4x^2 in the equation: -6x-1 + g(x)

OpenStudy (anonymous):

-6x-1+4x^2 like this?

OpenStudy (anonymous):

Yes! :))) <3

OpenStudy (anonymous):

That's the answer! :) Good job! <3

OpenStudy (anonymous):

Answer choice A I believe! Great work! :))

OpenStudy (anonymous):

tell me if I am wrong but these are the exact same or no?

OpenStudy (anonymous):

You're right! They are the exact same :)

OpenStudy (anonymous):

Ok, here we go. Do you know the basic equation of a parabola?

OpenStudy (anonymous):

heck no TBI

OpenStudy (anonymous):

was it wrong?

OpenStudy (anonymous):

well maybe this will stick a little for some time.

OpenStudy (anonymous):

equation of parabola = \[f(x) =a(x-h)^{2} +k\] where vertex = (h,k)

OpenStudy (anonymous):

what grade are you in may I ask?

OpenStudy (anonymous):

college... I have a brain injury

OpenStudy (anonymous):

we need to rewrite the equation -3x^2 + 7x + 7 = f(x) (-3x^2 + 7x ) +7 = f(x)

OpenStudy (anonymous):

well did you guess the answer or work it out? It's right though :)

OpenStudy (anonymous):

Aw that sucks... how did that even happen? That's horrible, man! I hope you feel better.

OpenStudy (anonymous):

its an injury that will never go away it lintels memory loss and it causes things to be harder to learn I got it in the Marine Corps